40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Sep 8 00:32:20 2017, awade, Summary, FSS, Schematic block diagram FSS loops CTN_FSS_south_blockdiagram.pdf
    Reply  Thu Sep 14 21:54:26 2017, awade, Summary, FSS, FSS modifications: series resistance (R1 and R2) to monitor points on servo board D040105 
       Reply  Mon Sep 18 00:46:08 2017, rana, Summary, FSS, FSS modifications: series resistance (R1 and R2) to monitor points on servo board D040105 
          Reply  Mon Sep 18 13:34:13 2017, awade, Summary, FSS, FSS modifications: series resistance (R1 and R2) to monitor points on servo board D040105 
             Reply  Tue Oct 10 00:49:34 2017, awade, Craig, Summary, FSS, Replacing U3 on north path (2010:005) FSS servo board. 20171009_220549_Ser6_ComPathTP1toTP5_09-10-2017_175407.pdffmoenfipkniealnd.png20171009_DebuggingNorthPathServoBoard.tar.gz
    Reply  Mon Oct 2 12:30:15 2017, rana, Summary, FSS, LISO & mfil 
       Reply  Wed Oct 18 11:19:15 2017, awade, Summary, FSS, Comparing south TF to LISO model TTFSS_schematics_awades_hand_notes.pdf001.pdf20171017_South_FSSTFs.tar.gz
          Reply  Wed Oct 18 15:42:59 2017, awade, Summary, FSS, Comparing south TF to LISO model 
             Reply  Mon Oct 23 15:05:44 2017, awade, Summary, FSS, Comparing south TF to LISO model Ser7_South_MainBoard_compiledTF.pdfSer5_North_MainBoard_compiledTF.pdf2017-10-20_18.04.18.jpg
                Reply  Wed Oct 25 23:15:51 2017, awade, Craig, Summary, FSS, Comparing south and north TF to LISO model South_MainBoard_EXC-30dBm_compiledTF.pdfNorth_MainBoard_EXC-30dBm_compiledTF.pdfFSS_all.tar.gz
Message ID: 1947     Entry time: Tue Oct 10 00:49:34 2017     In reply to: 1926
Author: awade, Craig 
Type: Summary 
Category: FSS 
Subject: Replacing U3 on north path (2010:005) FSS servo board. 

I had another look at the common path gain stage of the FSS field boxes (LIGO-D040105).  In modifications to the two field boxes in the PSL lab a switch with capacitor in series had been added to U7 for intergrator boost.  A gain killing switch, with 10 Ω in series, had also been added to U3 to manually kill gain while engauging the intergrator.  We suspect this -- along with only 100 Ω to ground on the out2 test point -- may have degrated the U3 AD829 chips. 

A transfer function from TP2 to TP5 of the north servo board (LIGO-D040105) was taken.  An excitation was put into the Test2 SMA (on RF board LIGO-D0901894) and a pair of HP 41800a active probes were used to take the TF accross the test points. Also, for comparison, I did the same for the FSS field box procured from the 40m (2010:006).  The data are plotted below: Ser5 is the north path serial numbered #5 board and Ser6 is the #6 board of the 40m unit.  Between the north path (2010:005) and 40m (2010:006) boxes there is a overall gain difference (due to changes in the feedback resistor) but the slope is very different.

I then replaced U3 (an AD829) on the servo board (LIGO-D040105) and retook the transfer function (also plotted below). With a new op amp the slopes of the magnitude and phase of the bode plots now match between (presumably working) 40m box and the north path FSS box. This is good.

One thing I don't understand is why there is a roll off starting at or below 1 kHz.  I used a 300 Hz resolution BW on the TF so below 600 Hz is bad.  But I've probably done something stupid in the way I've gone about this measurment.  I wanted something quick and easly repeatable, I'll probably realize tommorrow what I've done wrong.  Craig measured this about two weeks ago and it looks much flatter. What he sent me is attached below, it includes a LISO model. From the info on the AD829 datasheet we expected it to be flatter out to 1 MHz for a gain of 4.

So these seems replacing U3 on LIGO-D040105 board is a slight win but we (really just me) still don't understand why it has a sloped TF below 1 MHz.

---

I also took the opportunity to replace the R1 and R2 resistors with their design value of 453 Ω. I had used 470 Ω in a previous rework of the board (PSL:1918), but these were the wrong pitch and I did a messy job of it.  I found these 453 Ω taped to the inside of the north box.

 

Attachment 1: 20171009_220549_Ser6_ComPathTP1toTP5_09-10-2017_175407.pdf  131 kB  | Hide | Hide all
20171009_220549_Ser6_ComPathTP1toTP5_09-10-2017_175407.pdf
Attachment 2: fmoenfipkniealnd.png  64 kB  | Hide | Hide all
fmoenfipkniealnd.png
Attachment 3: 20171009_DebuggingNorthPathServoBoard.tar.gz  8 kB
ELOG V3.1.3-